Find three consecutive integers such that the sum of the first and twice the second is 122 minus three times the third.
Answer by checkley77(12844)

Let x, x+1 & x+2 be the 3 integers.
x+2(x+1)=122-3(x+2)
x+2x+2=122-3x-6
x+2x+3x=122-6-2
6x=114
x=19 ans. for the smaller integer.
19+1=20 ans. for the middle integer.
19+2=21 ans. for the largest integer integer.
Proof:
19+2*20=122-3*21
19+40=122-63
59=59